OpenStudy (anonymous):

the sum of 3 consecutive integers is -54. What is the largest integer?

OpenStudy (anonymous):

Let the first one be x. x + x + 1 + x + 2 = -54 Solve for x.

OpenStudy (anonymous):

Then, x + 2 will be the largest integer.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x + x + 1 + x + 2 = -54 }\) \(\Large \color{purple}{\rightarrow 3x + 3 = -54 }\) \(\Large \color{purple}{\rightarrow 3x = -517}\) \(\Large \color{purple}{\rightarrow x = -19 }\) So, the largest integer is x + 2 = -17

OpenStudy (anonymous):

x=-19